An atom has completely filled inned shells and a single valence electron in an excited p state. The filled inner shells have an orbital ommemtum eaual to zero. What is the magnitude of the orbital angular momentum of the atom?

A) 1.0 h/2pi

B) 1.2 h/2pi

C) 1.4 h/2pi

D) 1.7 h/2pi

E) 2.0 h/2pi

Can you please show the right answer and the work with detail? The professor says the answer is B but everyone in class got C.
Answer #1

for p statr angular momentum quantum number l = 1

orbital angular momentum L = sqrt(l(l+1)) h/2pi

L = sqrt(1(1+1))h/2pi

L = sqrt(2) h/2pi

L = 1.414 h/2pi

approximately L = 1.4 h/2pi